# suppReferences

### v4

@@ Introduction
Second, we removed tweets where Muse’s response references at least one related fact-checking article, accounting for 16% of all tweets. Notably, in only 6% of all tweets, Muse’s response exclusively references related fact-checking article(s). The quality of Muse-generated responses to these tweets can be more associated with Muse’s ability to summarize retrieved professional fact-checking articles, rather than fact-checking such content from scratch. We found that the quality of Muse-generated responses to content that has not been fact-checked online is at least 28% higher than responses by GPT-4 and even high-helpfulness responses by laypeople (Fig. 3b). The average quality score of Muse-generated responses for these potentially more challenging tweets is 8.0, close to its 8.1 average score for all tweets.
@@
Third, LLMs have exhibited political bias [47], which may diminish their ability to debunk misinformation that aligns with their political stance. We annotated the political leaning of tweets and found that GPT-4, which has an observed liberal bias [47], indeed performs 5% worse in correcting misinformation that aligns with liberal ideology compared to conservative ideology (Fig. 3c). Nevertheless, the quality of Muse-generated responses to both liberal and conservative content has the same average score of 8.3. This score is consistently higher than GPT-4-generated responses and even high-helpfulness responses from laypeople by at least 26%.
@@
Finally, we annotated the domain of tweets and tactics used to make them or part of them false or misleading. Results in Fig. 3d-e indicate that the quality of Muse-generated responses to potential misinformation across different domains and tactics maintains an average score around 8, consistently higher than responses by GPT-4 and even high-helpfulness responses from laypeople, with a margin of at least 19%. Here, domains broadly include politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine. Tactics primarily include fabricating, e.g., a story with or without an authentic image, or using a digitally fabricated or altered screenshot or photo; misinterpreting or misrepresenting someone’s claim, a symbol, a policy, and others; lacking context; implying false or oversimplified causation, such as conflating correlation with causation; using loaded language; presenting false, partial, or biased data, such as cherry-picking; and having improper analogy or equivalence (see examples in Supp. Fig. S7).
@@ Discussion
While concerns have arisen about LLMs in facilitating the creation of misinformation[48, 49], our work demonstrates LLMs’ potential to improve the online information ecosystem by correcting misinformation[2, 3]. Correcting misinformation requires identifying what part(s) of the content is (in)accurate and explaining why that part of the content (in)accurate with trustworthy references. Such identifications with clear explanations have been shown to reduce misinformation’s spread and people’s false beliefs, but its effectiveness is affected by quality, timeliness, and scalability[4, 22]. However, research on scaling up and accelerating misinformation correction is still at the early stage. Existing AI models are ill-equipped to identify and correct misinformation on social media. Such misinformation often spreads rapidly and is not restricted to narrow domains. It can combine (in)accurate, factually accurate but misleading, and unverifiable claims. It can employ various tactics, deliberately or accidentally, to create falsehood or misleadingness. Addressing it requires comprehending content that can be multimodal and the context beyond it that often involves emerging events. Existing methods to evaluate the quality of corrections presented in natural language also struggle to be accurate and comprehensive. Such evaluation relies on professional knowledge and skills. It should thoroughly assess the identification and explanation of (in)accuracies, as well as the use of references from diverse perspectives.
@@
We propose Muse, which augments existing powerful LLMs (here, it is GPT-4) with the capabilities of addressing images, accessing up-to-date knowledge, and finding accurate references, and 13 dimensions of misinformation correction quality. Evaluation by fact-checking experts demonstrates the high quality of Muse’s automatically generated responses to social media posts that potentially are misinformation. Results further validate that GPT-4 struggles to effectively respond to visual content but Muse excels (Fig. 3a). Muse also exhibits significantly fewer “hallucinations” by having fewer errors in identifying and explaining (in)accuracies (Fig. 2c,e), generating text that is more factual and relevant to the responded content (Fig. 2g-h), and providing more references that are real and relevant to the generated text (Fig. 2l-m) than GPT-4. Note that addressing LLM ’hallucination’ has been a long-standing challenge stressed in many works[26, 50].
@@
We provide Muse as a solution to assist social media users and platforms in reliably, scalably, promptly, and transparently responding to suspicious content. Muse is end-to-end and thereby simple to use. It is nonparametric and thereby easily and cheaply updated[51]. Meanwhile, our results reveal that Muse-generated responses have high quality in identifying and explaining inaccuracies, generated text, and provided references, significantly surpassing GPT-4 and laypeople who produce even high-helpfulness responses in correcting misinformation across modalities, domains, political leanings, and tactics and that has not even been fact-checked online (Fig. 2-3). Besides the highest accuracy and factuality (Fig. 2c,e,g-h,l-m), Muse’s generated responses show the highest readability by being the most explicit, fluent, and coherent (Fig. 2b,i-j). Muse can reduce the risk of a correction backfiring (i.e., reinforcing rather than reducing people’s false beliefs) by generating responses with the least toxic text and most credible references[29, 41, 42] (Fig. 2k,n). It can also reduce the implied truth effect (i.e., increasing people’s perceived accuracy of overlooked inaccuracies)[22] by comprehensively identifying all the inaccuracies in a social media post (Fig. 2d) and being more capable of correcting misinformation at scale. These high-quality responses by Muse can generally be obtained within minutes of suspicious content appearing on social media (Methods). By transparently providing references that refer to the retrieved web pages where evidence was collected, users can become more informed and also verify responses themselves (Fig 2,3f). Muse can also inform users of the current lack of evidence, provide an accuracy nudge, and when applicable, express uncertainty (Supp. Fig. S8).
@@
Muse’s responses cost about 0.5 USD per social media post at the time of our evaluation, though this cost has now been reduced to 0.2 USD, as GPT-4’s price has lowered (as of February 2024; Methods). Our focus in designing Muse was in maximizing the quality of corrections. Considering the task’s complexity and our significant improvement in quality, the cost is relatively inexpensive compared to alternatives. For example, a crowd of laypeople can already cost about 0.9 USD[19] to identify whether a new article’s headline and lede contains misinformation without writing down the explanation. Notably, we also changed Muse’s foundation LLM from GPT-4 to open-sourced Llama-3 (70B) and observed two Muses are comparable without modifying any other implementation details (Supp. Fig. S9; Methods), suggesting Muse’s generalizability. Note that the knowledge cutoff for Llama-3 is December 2023, whereas that of GPT-4 is September 2021. Llama-3’s training data might have included some or even all tweets used in evaluation and related events, knowledge, and corrections (e.g., their Community Notes data). As a comparison, GPT-4’s training data might only include up to 10% (Methods, where we show that Muse’s performance remains consistent after removing these tweets). Hence, it is unfair to compare between Llama-3-based and GPT-4-based Muse, and we did not include Llama-3 and its based Muse in expert evaluation. Nevertheless, our qualitative analysis illustrated in Supp. Fig. S9 reveals the potential to significantly reduce Muse’s cost, as GPT-4 was Muse’s most costly component. Furthermore, as one of the most capable open-sourced LLMs, Llama-3 still cannot accept visual inputs (as of August 2024), emphasizing the importance of Muse’s image describer module.
@@
This work also faces the following limitations. First, although Muse is capable of responding to multimodal misinformation with text and images, it cannot accept video inputs. Second, we only focus on English, one of the most spoken languages in the world. Third, we evaluated Muse using real social media content on a single platform, X, as its Community Notes system has been shown to reduce the spread of misinformation[17] and transparent. X is also a popular social media platform, and one where more than half of users consume news regularly[52] and where misinformation has been shown to diffuse faster than the truth[13]. Fourth, experts assessed and compared Muse against one foundation LLM, GPT-4, which can also be seen as an ablation study, since Muse augments GPT-4 (see Methods and Supp. Fig. S24 for additional results of our ablation study). GPT-4 was chosen as a comparison as it is one of the best performing LLMs across a wide range of tasks available today[53, 54].
Conclusion
We proposed Muse, a nonparametric LLM to identify and explain (in)accuracies in social media content. Muse can integrate textual and visual information, access timely knowledge, and generate responses that use natural languages and have accurate and trustworthy references. Experts assessed that Muse significantly outperforms GPT-4 and even laypeople who write high-helpfulness responses on X Community Notes in identifying and explaining (in)accuracies, generating high-quality text, providing high-quality references, and producing high-quality responses overall. Muse excels when the responded content is textual or multimodal, related to a broad range of domains, false or misleading by applying various tactics, and even when the content has not been fact-checked online. Muse also performs consistently well when responding to liberal and conservative content. This work demonstrates the potential of LLMs to respond to online misinformation in a scalable, prompt, reliable, and transparent manner—addressing key bottlenecks where existing studies struggle. This work also contributes to a comprehensive framework, incorporating with 13 associated standards (Methods, where we analyzed their importance), for evaluating misinformation correction models. Additionally, we publicly release rich annotations on social media posts covering aspects such as political leanings, domains, and tactics that make them false or misleading, along with expert-annotated responses to enable research.
@@ Acknowledgements
X.Z., A.X.Z., and T.A. designed the study. X.Z. led, and A.S. assisted in developing the model. A.S. built the web interface for evaluation. X.Z. prepared and analyzed the data. X.Z. drafted and all authors revised the manuscript. A.X.Z. and T.A. supervised the study.
Competing interests
The authors declare no competing interests.
Additional information
Supplementary information is available for this paper.
Correspondence and requests for materials should be addressed to Tim Althoff (althoff@cs.washington.edu) and Amy X. Zhang (axz@cs.uw.edu).
@@
Figure S7: Examples of social media content that uses various tactics to make it or part of it false or misleading. These tactics include a: Fabricating a news event (#1), a story with an authentic image (#2, also related to conspiracy theory), and an image with AI (#3). b: Lacking context regarding the motivation behind behavior. Here, it is the consideration that the bill contained no funding for Florida, without which could mislead readers. c: Misinterpreting or misrepresenting a plan, where Biden did not promise to cancel all the debt (#1), or symbol, which is not Nazi SS runes but a shorthand for the 46th Separate Airmobile Brigade (#2). d: Using loaded language. e: Improperly analogizing or equating masks blocking respiratory droplets and underwear blocking gas molecules (#1), and COVID-19 and cancer (#2). f: Presenting false, partial, or biased data. 365348 includes multiple reports for the same child and runaways who may not be considered missing because of their guardians know their whereabouts (#1). The data ends in 1885 and is from a specific high elevation site in Greenland, not global temperatures (#2). g: Implying false or oversimplified causation. Many COVID-19 deaths occurred among Medicare beneficiaries because Medicare primarily serves the groups who are at greater risk of adverse outcomes from COVID-19 (#1). The boy was shot when—but not because—he was campaigning for Raphael Warnock (#2). Note that one content may apply more than one tactic intentionally or accidentally.
@@
1Describe an image in an informative way. Your description should be only based on the given {short caption}, {name of each person}, and {raw text}. If the image is from social media, you should start with "A screenshot of". If the image is a quote from someone, you should start with "A quote from" followed by this person’s name if there is any, then by the quoted text. If the image is an article, you should start with "An article". If the image is a photo, you should start with "A photo of". If the image is a map, you should start with "A map of". {raw text} may contain nonsense data that are unnecessarily included in the image description; however, {name of each person} is not, and if the concept in {raw text} has a conflict with that in {short caption} (e.g., "Robbie Lemos" versus "robbie leems" shown later), {raw text} is often the right one.2short caption: {a woman with glasses and a quote that says, in real life, i assure you there is no such thing as algebra}3name of each person: {Fran Lebowitz}4raw text: {"In real life, I assure you, there is no such thing as algebra."}5image description: {A quote from Fran Lebowitz, "In real life, I assure you, there is no such thing as algebra."}6short caption: {two men in suits}7name of each person: {Jim Caviezel, Michael Emerson}8raw text: {}9image description: {A photo of Jim Caviezel and Michael Emerson in suits}10short caption: {robbie leems on twitter}11name of each person: {}12raw text: {Robbie Lemos @RobbieLemos 1d I’d like to congratulate my dear friend Deep Mind on a wonderful 1st day at work today at Google. Just in time for #EarthDay2023, cheers brother! 1 2 3,790}13image description: {A screenshot of a post of Robbie Lemos, "I’d like to congratulate my dear friend Deep Mind on a wonderful 1st day at work today at Google. Just in time for #EarthDay2023, cheers brother!" The post was posted on Twitter.}14short caption: {a moose}15name of each person: {}16raw text: {Yahoo Finance @YahooFinance Typically, the stock market bottoms four to five months before a recession ends, but RBC’s research details that it has bottomed as early as nine months before the end of a recession. finance.yahoo.com Could the stock market power through a recession? ’This would be rare.’ 09:57 22/4/2023 3.4,011 Views 1 Retweet 1 Quote 5 Likes 1 Bookmark}17image description: {A screenshot of a post from Yahoo Finance, "Typically, the stock market bottoms four to five months before a recession ends, but RBC’s research details that it has bottomed as early as nine months before the end of a recession." The post shared an article from finance.yahoo.com claiming, "Could the stock market power through a recession? ’This would be rare.’" with a picture of a moose. The post was posted at 09:57 22/4/2023.}18short caption: {a person pouring tea into a cup}19name of each person: {}20raw text: {New research reveals how coffee and tea can affect risk of early death for adults with diabetes By Sandee LaMotte, CNN Updated 7:01 PM EDT, Wed April 19, 2023 f The health benefits of tea 01:10 - Source: CNN}21image description: {An article claiming, "New research reveals how coffee and tea can affect risk of early death for adults with diabetes." It attached a picture of a person pouring tea into a cup. It was written by Sandee LaMotte, published by CNN, and updated at 7:01 PM EDT, Wed April 19, 2023.}22short caption: {two people standing next to each other with the words love is blind}23name of each person: {Nick Lachey}24raw text: {\"Love Is Blind\" co-host faceplants with a regressive line of questioning Hayley Miller MSNBC DAILY MSNBC}25image description: {An article claiming, "’Love Is Blind’ co-host faceplants with a regressive line of questioning." It attached a picture of Nick Lachey and another person standing next to each other. It was written by Hayley Miller and published by MSNBC.}26short caption: {a bar graph that shows how engaged are the most followed journalists on twitter}27name of each person: {Rahul Kanwal}28raw text: {How engaged are the most-followed journalists on Twitter? Percentage of tweets from each journalist that are at-replies BDUTT 64%29image description: {A bar graph showing how engaged the most followed journalists, including Rahul Kanwal, are on Twitter through the percentage of tweets from each journalist that are at-replies. The chart was made by mattmaldre.com.}30short caption: {a graph showing the global defense budget by region}31name of each person: {}32raw text: {Global Defense Budgets by Region (\$ Billions) \$1,000 800 600 400 200 0 2020 2021 2022 2023 2024 2025 Asia-Pacific Latin America North America Sub-Saharan Africa Europe Middle East & North Africa Russia & Commonwealth of Independent States Source: Aviation Week}33image description: {A graph showing the global defense budget by region. It is from Aviation Week.}34short caption: {[IMAGE_CAPTION]}35name of each person: {[CELEBRITIES]}36raw text: {[OCR]}37image description:
@@
Figure S12: Distribution of helpfulness scores of laypeople’s responses in X Community Notes. All: All laypeople’s responses in Community Notes. Average helpfulness: Laypeople’s responses in Community Notes identified with average helpfulness and used in our study. High helpfulness: Laypeople’s responses in Community Notes identified with high helpfulness and used in our study. For $x\pm y$, $x$: mean, $y$: standard deviation. Community Notes data are regularly updated; ours are up until February 12, 2023.
@@
(a) Correlation for all responses.
(b) Correlation for responses whose overall quality is relatively low vs high.
Figure S25: Spearman correlation coefficient between overall response quality and each criterion that specifies overall response quality.

### v5

@@ Introduction
Second, we removed posts where Muse’s response references at least one related fact-checking article, accounting for 16% of all posts. Notably, in only 6% of all posts, Muse’s response exclusively references related fact-checking article(s). The quality of Muse-generated responses to these posts can be more associated with Muse’s ability to summarize retrieved professional fact-checking articles, rather than fact-checking such content from scratch. We found that the quality of Muse-generated responses to content that has not been fact-checked online is at least 28% higher than GPT-4 and even high-helpfulness responses by laypeople (Fig. 3b). The average quality score of Muse’s responses for these potentially more challenging posts is 8.0, close to its 8.1 average score for all posts.
@@
Third, LLMs have exhibited political bias [45], which may diminish their ability to debunk misinformation that aligns with their political stance. We annotated the political leaning of posts and found that GPT-4, which has an observed liberal bias [45], indeed performs 5% worse in correcting misinformation that aligns with liberal ideology compared to conservative ideology (Fig. 3c). Nevertheless, the quality of Muse-generated responses to both liberal and conservative content has the same average score of 8.3. This score is consistently higher than GPT-4 and even high-helpfulness responses from laypeople by at least 26%.
@@
Finally, we annotated the domain of posts and tactics used to make them or part of them false or misleading. Results in Fig. 3d-e indicate that the quality of Muse’ responses to potential misinformation across different domains and tactics maintains an average score around 8, consistently higher than responses by GPT-4 and even high-helpfulness responses from laypeople, with a margin of at least 19%. Here, domains broadly include politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine. Tactics primarily include fabricating, e.g., a story with or without an authentic image, or using a digitally fabricated or altered screenshot or photo; misinterpreting or misrepresenting someone’s claim, a symbol, a policy, and others; lacking context; implying false or oversimplified causation, such as conflating correlation with causation; using loaded language; presenting false, partial, or biased data, such as cherry-picking; and having improper analogy or equivalence (see examples in Supp. Fig. S7).
We further investigated the impact of potential training data contamination and poster popularity, demonstrating that Muse’s superior performance generalizes to correcting misinformation not included in its training data, as well as misinformation posted by users with varying levels of popularity (Methods). We also conducted a detailed ablation study, showing the effectiveness of the proposed retrieval mechanism and multimodal integration (Methods). Beyond expert evaluation, we conducted an additional study with 988 laypeople to examine how social media users perceive and react to corrections generated by Muse (Methods). The results demonstrate Muse’s superior performance in generating corrections useful to end users. Notably, it is the only approach shown to significantly increase people’s capability to correctly identify misinformation, with an improvement of 9.8%, compared to even strong baselines, including GPT-4 and high-quality responses from laypeople.
@@ Discussion
While concerns have arisen about LLMs in facilitating the creation of misinformation[46, 47], our work demonstrates LLMs’ potential to improve the online information ecosystem by correcting misinformation[11, 3]. Correcting misinformation requires identifying which part(s) of the content are (in)accurate and explaining why with trustworthy references. This approach can reduce misinformation sharing and false beliefs, but effectiveness depends on the correction’s quality, timeliness, and scalability [4, 21]. Existing approaches, still at an early stage, struggle in particular with complex, nuanced misinformation that spans domains and goes beyond straightforward factual errors, especially when addressing it requires comprehending multimodal content and situating it in rapidly evolving real-world context. Existing methods for evaluating the quality of natural-language corrections also struggle to be accurate and comprehensive. Such evaluation relies on rigorous assessment of the identification and explanation of (in)accuracies, as well as the use of references from diverse perspectives. Beyond content validity, evaluation should also probe likely real-world impact by examining how end users perceive and act on the correction.
@@
We propose Muse, which augments existing powerful LLMs (here, it is GPT-4) with the capabilities of addressing images, accessing up-to-date knowledge, and finding accurate references, as well as 13 dimensions of correction quality. Evaluation by fact-checking experts demonstrates the high quality of Muse-generated responses to social media posts as potential misinformation. Results further validate that GPT-4 struggles to effectively respond to visual content but Muse excels (Fig. 3a). Muse also exhibits significantly fewer “hallucinations” by having fewer errors in identifying and explaining (in)accuracies (Fig. 2c,e), generating text that is more factual and relevant to the responded content (Fig. 2g-h), and providing more references that are real and relevant to the generated text (Fig. 2l-m) than GPT-4. Note that addressing LLM ’hallucination’ has been a long-standing challenge stressed in many works[22, 48].
@@
As more and more social media platforms (e.g., Facebook, Instagram, Threads [12]) are adopting Community Notes-like approaches to address misinformation, we provide Muse as a potential solution to assist social media users in drafting responses and to help platforms directly respond to misinformation reliably, promptly, and transparently. Muse is end-to-end and thereby simple to use. It is nonparametric and thereby easily and cheaply updated[49]. Meanwhile, our results reveal that Muse-generated responses have high quality in identifying and explaining inaccuracies, generated text, and provided references, significantly surpassing GPT-4 and laypeople who produce even high-helpfulness responses in correcting misinformation across modalities, domains, political leanings, and tactics and that has not even been fact-checked online (Fig. 2-3). Besides the highest accuracy and factuality (Fig. 2c,e,g-h,l-m), Muse’s generated responses show the highest readability by being the most explicit, fluent, and coherent (Fig. 2b,i-j). Muse can reduce the risk of a correction backfiring (i.e., reinforcing rather than reducing people’s false beliefs) by generating responses with the least toxic text and most credible references[25, 37, 38] (Fig. 2k,n). It can also reduce the implied truth effect (i.e., increasing people’s perceived accuracy of overlooked inaccuracies)[21] by comprehensively identifying all the inaccuracies in a social media post (Fig. 2d) and being more capable of correcting misinformation at scale. Additionally, we empirically demonstrate that Muse-generated responses enhance people’ capability to correctly identify misinformation by 9.8% (Methods). These high-quality responses and effective corrections by Muse can generally be obtained within minutes of suspicious content appearing on social media (Methods). By transparently providing references that refer to the retrieved web pages where evidence was collected, users can become more informed and also verify responses themselves (Fig 2,3f). Muse can also inform users of the current lack of evidence, provide an accuracy nudge, and when applicable, express uncertainty (Supp. Fig. S8).
@@
Muse’s responses cost about 0.5 USD per social media post at the time of our evaluation, though this cost has now been reduced to 0.2 USD, as GPT-4’s price has lowered (as of February 2024; Methods). Our focus in designing Muse was in maximizing the quality of corrections. Considering the task’s complexity and our significant improvement in quality, the cost is relatively inexpensive compared to alternatives. For example, a crowd of laypeople can already cost about 0.9 USD[18] to identify whether a new article’s headline and lede contains misinformation without writing down the explanation. We also changed Muse’s foundation LLM from GPT-4 to open-sourced Llama-3 (70B) and observed two Muses are comparable without modifying any other implementation details (Supp. Fig. S9; Methods), suggesting Muse’s generalizability. Note that the knowledge cutoff for Llama-3 is December 2023, whereas that of GPT-4 is September 2021. Llama-3’s training data might have included some or even all posts used in evaluation and related events, knowledge, and corrections (e.g., their Community Notes data). As a comparison, GPT-4’s training data might only include up to 10% (Methods, where we show that Muse’s performance remains consistent after removing these posts). Hence, it is unfair to directly compare between Llama-3-based and GPT-4-based Muse, and we did not include Llama-3 and its based Muse in evaluation. Nevertheless, our qualitative analysis illustrated in Supp. Fig. S9 reveals the potential to significantly reduce Muse’s cost, as GPT-4 was Muse’s most costly component. Furthermore, as one of the most capable open-sourced LLMs, Llama-3 still cannot accept visual inputs (as of August 2024), emphasizing the importance of Muse’s informative image captioning module.
@@
This work faces several limitations. First, although Muse is capable of responding to misinformation with text and images, it cannot accept video inputs. Second, we only focus on English, one of the most spoken languages in the world. Third, we evaluated Muse using real social media content on a single platform, X, as its Community Notes system has been shown to reduce the spread of misinformation[16] and transparent. X is also a popular social media platform, and one where more than half of users consume news regularly[50] and where misinformation has been shown to diffuse faster than the truth[12]. Fourth, experts assessed and compared Muse against one foundation LLM, GPT-4, which can also be seen as an ablation study, since Muse augments GPT-4 (see Methods and Supp. Fig. S24 for additional results of our ablation study). GPT-4 was chosen as a comparison as it is the state-of-the-art ChatGPT (as of April 2025) and one of the best performing LLMs across a wide range of tasks available today[51, 52].
@@ Acknowledgements
Supplementary material
Supplementary material is available for this paper.
@@
X.Z., A.X.Z., and T.A. conceptualized the work and developed the overarching research goals and aims. X.Z. led the model development. A.S. led the assessment of end-user perceptions of corrections, and X.Z. led the remainder of the evaluation. X.Z. drafted the manuscript, including visualization, and all authors revised the paper. A.X.Z. and T.A. supervised the study.
Data availability
Data, including code, used in this study are available at https://github.com/Social-Futures-Lab/MUSE. We comply with X Terms of Service by only releasing the IDs of tweets. The experts’ names are anonymized.
@@
Figure S7: Examples of social media content that uses various tactics to make it or part of it false or misleading. These tactics include a: Fabricating a news event (#1), a story with an authentic image (#2, also related to conspiracy theory), and an image with AI (#3). b: Lacking context regarding the motivation behind behavior. Here, it is the consideration that the bill contained no funding for Florida, without which could mislead readers. c: Misinterpreting or misrepresenting a plan, where Biden did not promise to cancel all the debt (#1), or symbol, which is not Nazi $S$ runes but a shorthand for the 46th Separate Airmobile Brigade (#2). d: Using loaded language. e: Improperly analogizing or equating masks blocking respiratory droplets and underwear blocking gas molecules (#1), and COVID-19 and cancer (#2). f: Presenting false, partial, or biased data. 365348 includes multiple reports for the same child and runaways who may not be considered missing because of their guardians know their whereabouts (#1). The data ends in 1885 and is from a specific high elevation site in Greenland, not global temperatures (#2). g: Implying false or oversimplified causation. Many COVID-19 deaths occurred among Medicare beneficiaries because Medicare primarily serves the groups who are at greater risk of adverse outcomes from COVID-19 (#1). The boy was shot when—but not because—he was campaigning for Raphael Warnock (#2). Note that one content may apply more than one tactic intentionally or accidentally.
@@
   1   {internallinenumbers*} Describe an image in an informative way. Your description should be only based on the given {short caption}, {name of each person}, and {raw text}. If the image is from social media, you should start with "A screenshot of". If the image is a quote from someone, you should start with "A quote from" followed by this person’s name if there is any, then by the quoted text. If the image is an article, you should start with "An article". If the image is a photo, you should start with "A photo of". If the image is a map, you should start with "A map of". {raw text} may contain nonsense data that are unnecessarily included in the image description; however, {name of each person} is not, and if the concept in {raw text} has a conflict with that in {short caption} (e.g., "Robbie Lemos" versus "robbie leems" shown later), {raw text} is often the right one.    2   {internallinenumbers*} short caption: {a woman with glasses and a quote that says, in real life, i assure you there is no such thing as algebra}    3   name of each person: {Fran Lebowitz}    4   raw text: {"In real life, I assure you, there is no such thing as algebra."}    5   image description: {A quote from Fran Lebowitz, "In real life, I assure you, there is no such thing as algebra."}    6   short caption: {two men in suits}    7   name of each person: {Jim Caviezel, Michael Emerson}    8   raw text: {}    9   image description: {A photo of Jim Caviezel and Michael Emerson in suits}    10   short caption: {robbie leems on twitter}    11   name of each person: {}    12   {internallinenumbers*} raw text: {Robbie Lemos @RobbieLemos 1d I’d like to congratulate my dear friend Deep Mind on a wonderful 1st day at work today at Google. Just in time for #EarthDay2023, cheers brother! 1 2 3,790}    13   {internallinenumbers*} image description: {A screenshot of a post of Robbie Lemos, "I’d like to congratulate my dear friend Deep Mind on a wonderful 1st day at work today at Google. Just in time for #EarthDay2023, cheers brother!" The post was posted on Twitter.}    14   short caption: {a moose}    15   name of each person: {}    16   {internallinenumbers*} raw text: {Yahoo Finance @YahooFinance Typically, the stock market bottoms four to five months before a recession ends, but RBC’s research details that it has bottomed as early as nine months before the end of a recession. finance.yahoo.com Could the stock market power through a recession? ’This would be rare.’ 09:57 22/4/2023 3.4,011 Views 1 Retweet 1 Quote 5 Likes 1 Bookmark}    17   {internallinenumbers*} image description: {A screenshot of a post from Yahoo Finance, "Typically, the stock market bottoms four to five months before a recession ends, but RBC’s research details that it has bottomed as early as nine months before the end of a recession." The post shared an article from finance.yahoo.com claiming, "Could the stock market power through a recession? ’This would be rare.’" with a picture of a moose. The post was posted at 09:57 22/4/2023.}    18   short caption: {a person pouring tea into a cup}    19   name of each person: {}    20   {internallinenumbers*} raw text: {New research reveals how coffee and tea can affect risk of early death for adults with diabetes By Sandee LaMotte, CNN Updated 7:01 PM EDT, Wed April 19, 2023 f The health benefits of tea 01:10 - Source: CNN}    21   {internallinenumbers*} image description: {An article claiming, "New research reveals how coffee and tea can affect risk of early death for adults with diabetes." It attached a picture of a person pouring tea into a cup. It was written by Sandee LaMotte, published by CNN, and updated at 7:01 PM EDT, Wed April 19, 2023.}    22   short caption: {two people standing next to each other with the words love is blind}    23   name of each person: {Nick Lachey}    24   {internallinenumbers*} raw text: {\"Love Is Blind\" co-host faceplants with a regressive line of questioning Hayley Miller MSNBC DAILY MSNBC}    25   {internallinenumbers*} image description: {An article claiming, "’Love Is Blind’ co-host faceplants with a regressive line of questioning." It attached a picture of Nick Lachey and another person standing next to each other. It was written by Hayley Miller and published by MSNBC.}    26   short caption: {a bar graph that shows how engaged are the most followed journalists on twitter}    27   name of each person: {Rahul Kanwal}    28   {internallinenumbers*} raw text: {How engaged are the most-followed journalists on Twitter? Percentage of tweets from each journalist that are at-replies BDUTT 64% RealMikeWilbon 38% 34% ErinAndrews 31% stephenasmith 24% andersoncooper rahulkanwal 13% 13% ninagarcia 5% sardesairajdeep 1% BillSimmons maddow 1% 0% 10% 20% 30% 40% 50% 60% chart by mattmaldre.com}    29   {internallinenumbers*} image description: {A bar graph showing how engaged the most followed journalists, including Rahul Kanwal, are on Twitter through the percentage of tweets from each journalist that are at-replies. The chart was made by mattmaldre.com.}    30   short caption: {a graph showing the global defense budget by region}    31   name of each person: {}    32   {internallinenumbers*} raw text: {Global Defense Budgets by Region (\$ Billions) \$1,000 800 600 400 200 0 2020 2021 2022 2023 2024 2025 Asia-Pacific Latin America North America Sub-Saharan Africa Europe Middle East & North Africa Russia & Commonwealth of Independent States Source: Aviation Week}    33   image description: {A graph showing the global defense budget by region. It is from Aviation Week.}    34   short caption: {[IMAGE_CAPTION]}    35   name of each person: {[CELEBRITIES]}    36   raw text: {[OCR]}    37   image description:
@@
Figure S12: Distribution of helpfulness scores of laypeople’s responses in X Community Notes. All: All laypeople’s responses in Community Notes. Average helpfulness: Laypeople’s responses in Community Notes identified with average helpfulness and used in our study. High helpfulness: Laypeople’s responses in Community Notes identified with high helpfulness and used in our study. For $\pm y, xx$: mean, $y$: standard deviation. Community Notes data are regularly updated; ours are up until February 12, 2023.
@@
Figure S25: Impact of popularity of users who posted tweets on Muse’s performance.
Figure S26: Study interface to measure end user perceptions of corrections (page 1/2, continued on the next page).
Figure S27: Study interface to measure end user perceptions of corrections (page 2/2).
Figure S28: We conducted a study on Prolific with 988 participants to assess the real-world impact of Muse and how people perceive and react to AI-generated responses (two-sided Student’s t-test; experiments=16). (a) Only Muse-generated corrections significantly enhanced the correct belief that the misinformation is misleading ($p=0.0001$). (b) None of the corrections made by laypeople, GPT-4, and Muse significantly impacted participants’ intention to share the misinformation, and participants’ intentions were generally low (around 2/7). (c) Muse, GPT-4, and high-helpfulness responses from laypeople all received significantly higher trustworthiness ratings than average-helpfulness responses from laypeople. (d-f) While disclosing the source of the correction (human or AI) had no significant effect on trustworthiness or intention to share, not disclosing the source significantly enhanced the correct belief that the misinformation is misleading ($p=0.0001$).